A vessel is filled with liquid,  3 parts of which are water and 5 parts syrup. How much of the mixture must be drawn off and replaced with water so that the mixture may be half water and half syrup.
Logically how it is to be done.  Thanks

Grade:8

1 Answers

Harsh Patodia IIT Roorkee
8 years ago
Water / Syrup = 3x/5x
Let as say y litres is taken out and replaced. Water taken out 3y/8 and syrup 5y/8
Acc to question

(3x- 3y/8 +y )/(5x-5y/8) = 1/1 (Since both are equal)
3x+5y/8 = 5x-5y/8
2x=5y/4
y = 8x/5
Total mixture 8x
Part that should be taken out is y/x = 1/5 th of original mixture.